A regular polygon has 14 sides what is the degrees of each interior angle?
Answer by Cromlix(4381)

Sum of the interior angles in a polygon
is given by: (n - 2) * 180
n = 14 (14 - 2) = 180
= 2160
2160/14 = 154.3 degrees (each interior angle)